OpenStudy (anonymous):

You buy a house including the land it sits on for $ 158000. The real estate agent tells you that the land costs $ 18000 more than the house. How much is the House? How much is the Land?

OpenStudy (radar):

Let x = cost of house land cost=x+$18000 ($18000 ore than house) $158000 =x+x+$18000=$158000. cost of house including land (given) $158000=2x+$18000 2x=$158000-$18000=$140000 Then x=$140000/2=$70000 (house cost) Land cost=$70000+$18000=$88000

OpenStudy (anonymous):

Thanks! that makes sense!